OpenStudy (anonymous):

The numerator of a fraction is 8 times the denominator. If 5 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 11/2. Find the original fraction.

OpenStudy (anonymous):

\[ n=8d\\ \frac{n-5}{d+5}=\frac{11}{2} \]

OpenStudy (anonymous):

So we can multiply both sides of the second equation by \(d+5\), to get \(n-5=\frac{11}{2}(d+5)\). So then \(n=\frac{11}2d+\frac{55}2+5,\ 2n=11d+60,\ 16d=11d+60,\ 5d=60\)... can you solve it from there?

OpenStudy (anonymous):

This method might be ahead of what you've learned so far.

OpenStudy (anonymous):

...I don't recognize it so that seems most likely 'x3

OpenStudy (anonymous):

What method did you use to solve problems like this in class?

OpenStudy (anonymous):

Online course with glitchy slideshows that don't explain things well... . _ .''

OpenStudy (anonymous):

What course is it? Which grade, or what math level?

OpenStudy (anonymous):

I'm in 12th but I failed part of Algebra freshman year and the just told me like...a month ago that I needed to make it up ~ _ ~''

OpenStudy (anonymous):

Okay, this is something you should know then. This is basic algebra. You make a variable to represent the numerator, which I called \(n\), and a variable to represent the denominator, which I called \(d\), and then you write the word problem out using equations.